OpenStudy (mtalhahassan2):

In a percentage composition investigation a compound was decomposed into its elements: 20.0 g of calcium, 6.0 g of carbon, and 24.0 g of oxygen. Determine the percentage composition of this compound.

OpenStudy (haseeb96):

given mass divided by its atomic mass so for Ca = 20 /40 =1/2 Ca =0.5 for C = 6 /12 =1/2 =0.5 for O = 24/16 =3/2 =1.5 okay? have u got it